OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today?

OpenStudy (anonymous):

put x = number of one step equations, then 1150 - x equal number of two step equations (since their total is 1150 then solve \[3x-2(1150-x)=1300\]for x

OpenStudy (anonymous):

\[3x-2300+2x=1300\] \[5x-2300=1300\] \[5x=3600\] \[x=720\]

OpenStudy (anonymous):

what the monkey is a one step or two step equation?